there are three consecutive integers. half the sum of the second and third number is 35. what are the integers
Answer by checkley77(12844)

Let x, x+1 & x+2 be the three integers.
(x+1+x+2)/2=35
2x+3=70
2x=70-3
2x=67
x=67/2
x= 33.5 ans. for the smallest integer.
33.5+1=34.5 ans. for the second integer.
33.5+2=35.5 ans. for the third integer.
Proof:
(34.5+35.5)/2=35
70/2=35
35=35